OpenStudy (anonymous):

What is the equation of the line that passes through the point (-2, 2) and has a slope of 5? Thank you to anyone who can explain how to do this!

OpenStudy (wwe123):

1) slope = 2, point (1, 6) (y - 6) = 2(x - 1) y - 6 = 2x - 2 y = 2x + 4 2) slope = -1/5, point (5, 4) (y - 4) = (-1/5)(x - 5) y - 4 = (-1/5)x + 1 y = (-1/5)x + 5 3) slope = 4/3, point (12, 1) (y - 1) = (4/3)(x - 12) y - 1 = (4/3)x - 16 y = (4/3)x - 15 4) slope = 0, y-intercept -3 y = -3 5) slope = 2, point (-3, -10) (y + 10) = 2(x + 3) y + 10 = 2x + 6 y = 2x - 4 6) slope = 1/2, point = (-2, 4) (y - 4) = (1/2)(x + 2) y - 4 = (1/2)x + 1 y = (1/2)x + 5 7) slope = 5/4, point (4, -8) (y + 8) = (5/4)(x - 4) y + 8 = (5/4)x - 5 y = (5/4)x - 13 There, you have your lines

OpenStudy (azureilai):

use the point slope form.\[(y-y_{1})=m(x-x_{1})\] plug in (-2,2) for \[y_{1}\] and \[ x_{1}\] m = slope solve and isolate y to one side for y - intercept then rewrite equation in y=mx+b form

OpenStudy (anonymous):

you need to use the point slope formula because you have a set of points and a slope. y - y1 = m(x - x1) slope(m) = 5 (-2,2) x1 = -2 and y1 = 2 now just sub your info in the formula y - 2 = 5(x - (-2) y - 2 = 5(x + 2) y - 2 = 5x + 10 y = 5x + 10 + 2 y = 5x + 12

OpenStudy (anonymous):

Thank you for explaining. I originally had 2= -10 + 12. I was forgetting to add the two on the other side. :P

OpenStudy (anonymous):

easy mistake.....you would have eventually found your mistake and corrected it :)

OpenStudy (anonymous):

Thank you again for the help (: